### v2

# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

Abstract
  1. Accurate uncertainty quantification of large language models (LLMs) provides credibility measure over their outputs. However, fine-tuned LLMs often struggle with overconfidence in uncertain predictions due to the limitations in the models’ ability to generalize with limited data. Existing parameter efficient fine-tuning (PEFT) uncertainty quantification methods for LLMs focus on post fine-tuning stage and fall short of calibrating epistemic uncertainty. To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates epistemic uncertainty over the space of functions that map input prompts to outputs. We implement UQ4CT during the fine-tuning stage via a mixture-of-experts framework that hierarchically decomposes the functional space. We demonstrate that UQ4CT reduces Expected Calibration Error (ECE) by more than $25\%$ while maintaining high accuracy across $5$ benchmarks. Even under distribution shift, UQ4CT maintains superior ECE performance with high accuracy, showcasing improved generalizability.
11footnotetext: Department of Computer Science and Engineering, University of California San Diego, La Jolla, California, USA22footnotetext: Halıcıoğlu Data Science Institute, University of California San Diego, La Jolla, California, USAfootnotetext: Preprint. Under Review. Correspondence to: Yian Ma <yianma@ucsd.edu>.
1 Introduction
Quantifying credibility of the outputs has been one of the most important problems around large language models (LLMs) (Chang et al., 2024). In particular, fine-tuned LLMs often struggle with overconfidence in gauging the uncertain of their outputs—oftentimes due to limited training data—and does not reflect the true credibility of the answers that they generate (Xiao et al., 2022a; Zhou et al., 2022; He et al., 2023; Tian et al., 2023; OpenAI, 2023). Such overconfidence has become a crucial problem in safety-related decision making and in fields where data is relatively limited, such as formal proof generation, climate science, and healthcare (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022). Methods that enhance uncertainty quantification of fine-tuned LLMs are sought after for the assurance of trustworthy predictions in applications.
One salient challenge of uncertainty quantification in large models is the trade-off among accuracy, calibration, and tractability. One would first like to maintain the capability of the LLMs and not sacrifice accuracy of the generated outputs. Given this desideratum, a number of existing works focus on perturbing the prompts that are fed into the language models and quantify the uncertainty associated with such perturbations (Hou et al., 2023; Farquhar et al., 2024; Gao et al., 2024). These methods capture the aleatoric uncertainty assuming that the trained model corresponds to the ground-truth data distribution. However, these approaches cannot capture the epistemic uncertainty that arises from the model being trained on limited data. To capture this epistemic uncertainty, one would need to introduce changes or perturbations to the model in accordance with the data distribution. On the tractability side, it is not realistic to consider perturbations to all the model parameters for an LLM (Minaee et al., 2024).
We therefore propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT) to calibrate the epistemic uncertainty in LLMs. Since we only focus on the epistemic uncertainty over the functional space that maps input prompts to the outputs, we no longer need to consider uncertainty in the entire parameter space. Instead, we only need to measure the uncertainty regarding the functional relationships. Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the interesting functional relationships arising from the underlying data.
In particular, we make use of the low rank adaptation (LoRA, Hu et al. (2021a)) ensemble at each layer of the LLM to form the basis functions. We then utilize the mixture of experts (MoE) architecture (Li et al., 2024) to hierarchically combine the basis functions and form the functional space. During fine-tuning, UQ4CT simultaneously learns the parameters in the LoRA experts and the prompt-dependent mixture weights to output the calibrated distribution over the functional space. During inference, the LoRA experts capture different functional relationships in the fine-tuning data throughout training, and the MoE routers dynamically select these functional bases conditioned on the input. The selection process models the functional level epistemic uncertainty, and consequently captures the uncertainty in the output space. We calibrate functional level epistemic uncertainty to align with predictive correctness during training time. This approach significantly improves uncertainty estimations of the model on its predictions without compromising accuracy. To summarize, our contributions include:
  • A novel epistemic uncertainty quantification approach with MoE architecture during fine-tuning to quantify functional level epistemic uncertainty and align with predictive correctness, which mitigates overconfidence issue and improves generalizability.
  • A novel calibration loss function incorporating predictive correctness to calibrate the prompt-dependent LoRA mixture for better uncertainty estimation.
  • More than $25\%$ expected calibration error (ECE) reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task; improved ECE performance under distribution shifts on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks without compromising accuracy.
2 Preliminaries
2.1 Low-rank Adaptation (LoRA)
LLMs have numerous large weight matrices to perform matrix multiplication, denoted as $\mathbf{W}_{0}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{in}}}$ that maps inputs $\mathbf{x}$ to outputs $\mathbf{h}$. Hu et al. (2021a) proposes LoRA, which fixes $\mathbf{W}_{0}$ and introduces a low-rank perturbation $\Delta\mathbf{W}$ to the weight matrix:
$\mathbf{h}$ $=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}.$ (1)
Here, $\Delta\mathbf{W}$ is calculated as the product of two matrices, $\mathbf{B}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{lr}}}$ and $\mathbf{A}\in\mathbb{R}^{n_{\text{lr}}\times n_{\text{in}}}$ where $n_{\text{lr}}$ is significantly smaller than $n_{\text{in}}$ or $n_{\text{out}}$. For example, we use $n_{\text{lr}}=32$ while $n_{\text{in}}=n_{\text{out}}=4096$ for the Llama2-7b model (Touvron et al., 2023a). Therefore, the total number of LoRA parameters for this $\Delta\mathbf{W}$ is $n_{\text{lr}}(n_{\text{in}}+n_{\text{out}})$, which is far smaller than the parameter count of the full matrix, $n_{\text{in}}n_{\text{out}}$. One of the key motivations of incorporating LoRA to fine-tune LLMs is the vast amount of memory cost reduction compared with fine-tuning on the full model. For an LLM with $7$ billion parameters, maintaining the average gradient and average squared gradients for optimization multiplies the memory required by a factor of $3$ compared to simply loading model weights. LoRA greatly mitigates this memory cost as the tripled memory consumption only applies to LoRA adapters.
2.2 Mixture of Experts (MoE)
LoRA Mixture-of-Experts (Li et al., 2024; Wu et al., 2024a) is an efficient approach to scale the number of parameters while maintaining the same computational bounds. LoRA MoE utilizes the top-k router to assign each token to the LoRA experts (Lepikhin et al., 2020). The router is a linear layer that maps the input hidden state $\mathbf{h}$ to a probability distribution of candidate experts.
The plain transformer block in a large language model consists of the $q,k,v$ encoding layers ($\text{FFN}_{q,k,v}$), layer norm (LN) and feed forward layer (FFN), together with residual connections. Formally, given $\mathbf{h}_{o}=[s_{0},s_{1},...,s_{n}]$ as the tokenized input text, the output of $\ell$-th layer is generated as:
$\mathbf{z}^{\ell}=f_{attn}(\text{FFN}_{q,k,v}(\text{LN}(\mathbf{h}^{\ell-1})))+\mathbf{h}^{\ell-1},\quad\mathbf{h}^{\ell}=\text{FFN}(\text{LN}(\mathbf{z}^{\ell}))+\mathbf{z}^{\ell}.$ (2)
Here, $f_{attn}$ represents the attention calculation step.
Let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{1\times d}$ $(1\leq i\leq s,1\leq\ell\leq L)$ denote the output hidden state of the $i$-th token at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension. With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-n gate router $R(\cdot)$ chooses n experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-n}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (3)
Finally, we obtain the final MixLoRA prediction with:
$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}R^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}\cdot\mathbf{h}^{\ell}$ (4)
where $\mathbf{W}$ is the pretrained weights of the target fine-tuning layer and $\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}$ is the k-th LoRA expert.
With MixLoRA defined in Equation 4, we can apply MixLoRA layers at $q,k,v$ encoding and FFN layers:
$\mathbf{z}^{\ell}=f_{attn}(\text{MixLoRA}_{q,k,v}(\text{LN}(\mathbf{h}^{\ell-1})))+\mathbf{h}^{\ell-1},\quad\mathbf{h}^{\ell}=\text{MixLoRA}(\text{LN}(\mathbf{z}^{\ell}))+\mathbf{z}^{\ell}.$ (5)
Note that the calculation of $E^{\ell}_{k}$ involves $\mathbf{W}^{\ell}$, which represents the weight of the original ATTN or FFN layer.
2.3 Aleatoric and Epistemic Uncertainties
In machine learning models, uncertainty can be categorized into aleatoric (data-wise) and epistemic (model-wise) uncertainty (Hora, 1996; Hüllermeier and Waegeman, 2021). For LLMs, aleatoric uncertainty arises from the inherently ambiguous and context dependent nature of natural languages where a single phrase or sentence can have multiple valid interpretations in different contexts. Epistemic uncertainty is introduced by the model’s lack of knowledge due to limited learning capabilities, suboptimal modeling or sparse training data.
Epistemic uncertainty is highly related to several well-known limitations of generative models. For example, it has been observed that when an LLM is pretrained on a diverse range of text data, it is generally well-calibrated, i.e. the predicted probability of the next token generally aligns with what is observed in real text. However, after fine-tuning or alignment with human preferences, the calibration error deteriorates (Zhao et al., 2021; Achiam et al., 2023a). A related phenomenon is forgetting, where the performance of a fine-tuned LLM diminishes on tasks outside the scope of the target task (Lin et al., 2023; Luo et al., 2023).
Motivated by these observations, we explore functional-level epistemic uncertainty in generative models and aim to develop metrics that assess model performance on specific problem instances to fine-tune the parameter mixture.
3 Methodology
The high level goal of UQ4CT is to balance the exploration and exploitation of different LoRA experts during fine-tuning. In particular, we incorporate the functional-level epistemic uncertainty (FEU) to calibrate the prompt-dependent parameter mixture with LoRA MoE.
Assume that our answer $a$ is generated via a mixture of mechanisms or models $M$, conditioning on the input prompt $x$. Assume that $e(a)$ is an embedding of $a$ so that least squares distance is a natural distance on the space of $e(a)$. For an expressive enough model class $\mathcal{M}$, and a calibrated distribution $P(M|x)$ over the model class, we can measure the deviation of the generated answer to the “ideal” one $a_{*}=f_{*}(x)$ as:
$\mathbb{E}_{M\sim P(M|x)}\left[\mathbb{E}_{a\sim P(a|M,x)}\left[\|e(a)-e(a_{*})\|^{2}\right]\right]$
$=\mathbb{E}_{P(M|x)}\underbrace{\left[\mathbb{E}_{P(a|M,x)}\left[\|e(a)-\mathbb{E}_{P(a|M,x)}[e(a)]\|^{2}\right]\right]}_{\text{Aleatoric Uncertainty}}+\underbrace{\mathbb{E}_{P(M|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-e(a_{*})\|^{2}\right]}_{\text{Epistemic Uncertainty}}.$ (6)
Note that we hereby quantify uncertainty as a function of the input prompt $x$, since the distribution of model $M$ conditions on $x$. We hence name the task “functional-level uncertainty quantification”.
3.1 Functional-level Epistemic Uncertainty
Figure 1: Mixture of Experts (MoE) architecture to capture and calibrate functional-level epistemic uncertainty. Experts $B_{1...K}A_{1...K}$ capture functional relationships in the data throughout fine-tuning, the weights $\alpha_{1...K}$ quantify the uncertainty in selecting these functional bases conditioned on the input hidden state $h$, which is the semantic representation of the input token $x$. In the UQ4CT workflow, we align this uncertainty with predictive correctness. When the router makes a correct prediction, the loss reinforces this decision, thereby increasing the router confidence in its selection, which aligns with a lower epistemic uncertainty. Conversely, when the router makes an incorrect prediction, the loss penalizes this selection, potentially causing the router to distribute its probabilities more broadly across experts, which is indicative of higher epistemic uncertainty.
Motivated by the decomposition in Eq. (3) for least squares loss, we consider a general distance $d$, and formally define epistemic uncertainty as the variance caused by model training procedure. Specifically, we focus on the variation introduced in the model fine-tuning stage of LLMs. Mathematically, given prompt $x$, we consider the following definition of epistemic uncertainty:
$\text{Epistemic Uncertainty}=\mathbb{E}_{M\sim P(\cdot|x)}\mathbb{E}_{M^{\prime}\sim P(\cdot|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-\mathbb{E}_{P(a^{\prime}|M^{\prime},x)}[e(a^{\prime})]\|^{2})\right]$ (7)
Here, $a^{\prime}$ and $a$ represents the outputs sampled from the parameter mixtures $M^{\prime}$ and $M$ conditioned on the prompt $x$. The epistemic uncertainty measures the least squares distance between $e(a)$ and $e(a^{\prime})$.
3.2 Decomposition of the Functional Space
Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the functional relationships in the underlying data. Then the functional space can be decomposed into a mixture of the LoRA ensembles. Naively, this decomposition writes:
$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{1}=1}^{K}\sum_{k^{2}=1}^{K}\cdots\sum_{k^{L}=1}^{K}\alpha_{k^{1},k^{2},\dots,k^{L}}g_{k^{L}}^{L}\left(g_{k^{L-1}}^{L-1}\left(\dots g_{k^{1}}^{1}(\mathbf{x})\right)\right),$ (8)
$g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=E_{k^{\ell}}^{\ell}\left(f_{trans}^{\ell}(\mathbf{h}^{\ell})\right),$ (9)
where $f_{trans}^{\ell}$ represents the necessary non-fine-tuned operations within a transformer block (i.e. layer norm, attention calculation, etc.). This linear decomposition requires examining $K^{L}$ parameters to span the entire functional space.
We instead consider a hierarchical decomposition of $f(\mathbf{x})$:
$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{L}=1}^{K}\alpha^{L}_{k^{L}}g_{k^{L}}^{L}\Bigg{(}\sum_{k^{L-1}=1}^{K}\alpha^{L-1}_{k^{L-1}}g_{k^{L-1}}^{L-1}\Bigg{(}\dots\sum_{k^{1}=1}^{K}\alpha^{1}_{k^{1}}g_{k^{1}}^{1}(\mathbf{x})\Bigg{)}\Bigg{)}.$
Here the output function is determined by the combinations of the weights $\alpha^{\ell}_{k^{\ell}}$ across all the layers. As a result, we only need $K\cdot L$ parameters to represent various compositions of the functional relationships.
To further improve the efficiency of representation, we apply a sparse dynamic routing and choose the mixture weights $\alpha_{k^{\ell}}=R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}$ that depend on the input from the previous layer:
$\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}g^{\ell}_{k^{\ell}}(\mathbf{h}^{\ell}).$ (10)
3.3 Quantifying FEU with LoRA MoE Framework
We quantify the functional-level epistemic uncertainty (FEU) with the MoE architecture, which is represented by the embedding $e(a)$ in Eq. (7). As shown in Figure 1, the LoRA experts that are made of matrices $B^{1...L}_{1...K},A^{1...L}_{1...K}$ in Eq. (1) capture important functional relationships in the data during fine-tuning. We treat these functional relationships represented by the LoRA experts as basis functions. As shown in Eq. (10), conditional on the input prompts, the more complex functional relationships that recursively map inputs to outputs are represented as linear combinations (mixture of experts) of the basis functions, hierarchically through each layer.
Uncertainty quantification in this functional space reduces to quantifying the uncertainty of the weights over the basis functions. As shown in Figure 1, the weights $\alpha_{k}^{\ell}$ from the top-n router dynamically mix these basis functions given the input hidden state. The top-n weights that produce the final output hidden state quantify the functional level epistemic uncertainty in the function selection. With mixture of LoRA experts architecture, we follow routing mechanisms of the MoE layers as in Eq. (3) and (4). Specifically, we employ top-2 gate routers, which chooses $2$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$\vspace{-0.5mm}R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (11)
Given an input prompt $x$ with length $s$, we model functional-level epistemic uncertainty (FEU) by aggregating $R^{\ell}(\mathbf{h}^{\ell}_{i})$ over both layer dimension and sequence dimension:
$\text{FEU}(x)=\frac{1}{s}\sum^{s}_{i=1}\left[\frac{1}{L}\sum^{L}_{\ell=1}R^{\ell}(\mathbf{h}^{\ell}_{i})\right].$ (12)
3.4 Training Calibration Loss
The FEU model of the epistemic uncertainty enables us to calibrate the parameter mixtures against predictive accuracy for better aligned output distributions. Specifically for the MoE top-k routers, we design the following calibration loss for training:
$\mathcal{L}_{\mathrm{cal}}=\left\|\mathbbm{1}{\{\text{MixLoRA}(x)=y^{*}\}}-\text{FEU}(x)\right\|^{2},$ (13)
where the first term is an indicator function of whether the model prediction matches the ground truth $y^{*}$ given the prompt $x$. We employ a one-hot definition of the ground truth confidence. When the prediction from current mixture model matches the ground truth, the ground truth confidence is $1$. Otherwise, when the predictions do not match, the ground truth confidence is 00.
3.5 Calibrating the Mixture
Our proposed calibration loss $\mathcal{L}_{\mathrm{cal}}$ improves predictive reliability by adaptively balancing expert exploitation and exploration with functional-level epistemic uncertainty. As shown in Figure 1, $\mathcal{L}_{\mathrm{cal}}$ effectively promotes expert exploitation for correct predictions and expert exploration for incorrect predictions by directly calibrating the functional level epistemic uncertainty to align with the predictive correctness.
Ideally, when the $K$ LoRA experts together capture all the functional relationships across the data distribution with cross entropy during fine-tuning, our proposed loss $\mathcal{L}_{cal}$ also finds proper mixture of LoRA experts conditioned on the input $x$ by conditionally promoting expert exploitation and exploration. This allows the model to select correct functional relationships regarding $x$ to generate an output that better matches the data distribution, which grants calibrated uncertainty estimations.
Load balancing is a common regularization to ensure even utilization across experts with the MoE architecture (Fedus et al., 2022). We follow the load balancing loss $\mathcal{L}_{b}$ proposed by Li et al. (2024) and define our loss function as:
$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
where $\mathbf{CE}$ represents cross entropy loss, $\gamma$ and $\beta$ are the hyperparameters of two auxiliary terms. Details about $\mathcal{L}_{b}$ can be found in Appendix A.1.
4 Related Work
4.1 Parameter-Efficient Fine Tuning for LLMs
Large Language Models (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c) have shown impressive abilities in handling various natural language processing tasks. Building on these advances, instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) has enhanced LLMs’ capacity to comprehend and follow human instructions, forming the core of modern conversational AI systems (Wu et al., 2023b; Achiam et al., 2023b). However, as LLMs increase in size, the fine-tuning process demands much more time and memory.
To address these challenges, several strategies have been proposed, including parameter-efficient fine-tuning (PEFT) (Mangrulkar et al., 2022), model distillation (Liu et al., 2023; Xiao et al., 2023), quantization (Frantar et al., 2022; Xiao et al., 2022b), and pruning (Ma et al., 2023). Among these, LoRA (Hu et al., 2021b), which leverages low-rank matrix decomposition of linear layer weights, is a widely adopted PEFT technique that boosts model performance without adding computational costs during inference. For example, VeRA (Kopiczko et al., 2023) introduces learnable scaling vectors to modify shared pairs of frozen random matrices across layers, while FedPara (Hyeon-Woo et al., 2021) focuses on low-rank Hadamard products for federated learning settings. Tied-Lora (Renduchintala et al., 2023) applies weight tying to further minimize the number of trainable parameters. AdaLoRA (Zhang et al., 2023) uses Singular Value Decomposition (SVD) to prune less important singular values for efficient updates, and DoRA (Liu et al., 2024) separates pre-trained weights into magnitude and direction components, applying LoRA to update the directional component during fine-tuning, thus reducing the number of parameters to be trained.
4.2 Mixture of Experts
The Mixture-of-Experts concept (Jacobs et al., 1991; Wang et al., 2020), introduced as early as 1991, presented a novel supervised learning framework where multiple networks (experts) specialize in handling distinct subsets of training data. Modern MoE variants adapt this by modifying the traditional feed-forward sub-layer within transformer blocks, incorporating sparsely activated LoRA experts, which allows for significant expansion in model width without a proportional increase in computational overhead.
Different MoE architectures have since emerged, distinguished by their expert sampling and routing strategies. For example, LLaVA-MoLE (Chen et al., 2024) improves token routing to domain-specific experts within transformer layers, reducing data conflicts and consistently outperforming standard LoRA baselines. Other MoE-based approaches include MoRAL (Yang et al., 2024a), which focuses on efficiently adapting LLMs to new domains and tasks for lifelong learning, and LoRAMoE (Dou et al., 2024), which incorporates LoRAs via a router network to mitigate the issue of world knowledge forgetting. PESC (Wu et al., 2024b) transforms dense models into sparse ones through an MoE structure, lowering computational costs. MoE-LoRA (Luo et al., 2024) introduces a new parameter-efficient MoE method using Layer-wise Expert Allocation (MoLA) for transformer models, while MoCLE (Gou et al., 2023) activates task-specific model parameters based on instruction clusters. MixLoRA (Li et al., 2024) implements a high-throughput framework for LoRA MoE training and inference process, constructing LoRAs as stochastic experts to reduce computational overhead while expanding model capacity.
Despite the performance improvements these architecture advancements have brought, the overconfidence problem of fine-tuned models is lacking attention (Xiao et al., 2022a; He et al., 2023; Tian et al., 2023; OpenAI, 2023). Enhancing the uncertainty estimation capabilities of these models is fundamental toward more reliable, interpretable and trustworthy applications of LLMs.
4.3 Uncertainty Quantification in LLMs
Uncertainty quantification has garnered substantial attention in various tasks and domains within neural networks (Gal and Ghahramani, 2015, 2016a; Malinin and Gales, 2018; Ovadia et al., 2019; Dusenberry et al., 2020; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023). This focus extends to LLMs, where the precise quantification of prediction uncertainty has become a critical area of research (Xiao et al., 2022c; Lin et al., 2022; Mielke et al., 2022; Chen and Mueller, 2023; Duan et al., 2023; Huang et al., 2023). LLMs, particularly in generative tasks, pose unique challenges, especially when it comes to measuring the uncertainty of their outputs (Liu et al., 2019; Malinin and Gales, 2021; Kuhn et al., 2023; Lin et al., 2023). The distinction between aleatoric and epistemic uncertainty was recently examined in the context of LLMs (Hou et al., 2023), though this was approached by ensembling model inputs rather than model instances and did not address fine-tuning tasks specifically.
Established uncertainty quantification methods have been studied in conjunction with the LoRA structure. Monte-Carlo dropout (Gal and Ghahramani, 2016b) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them. Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty. Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023) combines the predictions from multiple LoRA adapters for better uncertainty calibration. Laplace-LoRA (Yang et al., 2024b) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates. However, none of these methods provide calibrations regarding epistemic uncertainty, which is crucial to mitigate overconfidence in the fine-tuning stage given the sparse dataset.
5 Experiments
5.1 Datasets
We include $5$ multiple-choice question answering benchmarks: OpenBookQA (OBQA, Mihaylov et al. (2018)), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) from AI2 Reasoning Challenge (Clark et al., 2018), BOOLQ (Clark et al., 2019) and ClimateQA, an expert-annotated domain specific benchmark for climate science. We also use computer science, law, medication and engineering subsets from MMLU dataset (Hendrycks et al., 2020) to evaluate performance under distribution shift. We fine-tune on the publicly available training split and test on the validation split from these benchmarks to evaluate model performance. We report the average model performance over $3$ random runs and the standard deviations in the subscript.
Table 1: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ LoRA $69.5_{1.93}$ $74.8_{1.39}$ $53.8_{0.6}$ $72.1_{0.87}$ $59.9_{2.13}$
MC Drop $66.8_{3.66}$ $76.8_{1.30}$ $50.9_{2.01}$ $74.8_{1.34}$ $58.2_{2.11}$
Ensemble $66.2_{3.7}$ $71.2_{1.0}$ $47.5_{0.57}$ $75.5_{1.4}$ $59.6_{6.9}$
LA $68.7_{1.32}$ $74.6_{2.11}$ $51.4_{0.83}$ $70.8_{1.24}$ $55.2_{3.29}$
MixLoRA $71.5_{1.05}$ $77.7_{2.27}$ $54.3_{1.07}$ $75.5_{2.91}$ $61.6_{1.76}$
UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
ECE $\downarrow$ LoRA $11.9_{0.78}$ $11.9_{2.04}$ $19.4_{4.75}$ $10.2_{1.07}$ $14.3_{0.64}$
MC Drop $12.2_{0.85}$ $11.9_{1.99}$ $19.8_{4.85}$ $10.9_{0.24}$ $14.3_{0.56}$
Ensemble $7.28_{2.3}$ $9.1_{1.49}$ $10.23_{1.39}$ $8.83_{2.35}$ $13.5_{3.29}$
LA $17.1_{1.72}$ $16.6_{3.7}$ $18.1_{0.5}$ $17.2_{1.2}$ $12.6_{1.9}$
MixLoRA $7.88_{2.09}$ $9.09_{0.81}$ $10.74_{1.07}$ $12.9_{1.99}$ $12.5_{1.32}$
UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
5.2 Experiment Setup
We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024) and compare the average performance in three random runs and report the mean and standard deviation with following baselines. We use the publicly available LLaMA-2-7B-hf (Touvron et al., 2023a) as our base model. In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-fowrward networks in LLaMA-2-7B-hf (gate layer, down layer and up layer). Details are provided in Appendix A.5
  • LoRA (Hu et al., 2021a). We use standard LoRA fine-tuning as lower performance bound.
  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016b) keeps dropout on at both training and testing time. By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions. To combine LoRA fine-tuning with MC dropout, we apply dropout on the input of the LoRA adapter, following the implementation of Mangrulkar et al. (2022).
  • Deep Ensemble (Lakshminarayanan et al., 2017) averages the predictions from each ensemble member which have been trained with varying random initialization. We combine deep ensemble with LoRA by fine-tuning $3$ randomly initialized LoRAs together and ensembling their output as final predictions.
  • Laplace-LoRA (LA) (Yang et al., 2024b) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
  • MixLoRA (Li et al., 2024) incorporates LoRAs via a router network to reduce computational overhead while expanding model capacity. We add this as a baseline to resemble plain LoRA MoE model performance.
Evaluation Metrics.
We measure the prediction accuracy on the validation set for all $5$ tasks. For uncertainty calibration, we incorporate expected calibration error (ECE, Guo et al. (2017)) with $15$ bins, which measures the alignment between predicted probabilities and empirical accuracy. We also investigate model performance under distribution shift to ensure the model has predictable behavior when given data from other domains as this is a crucial component for real-world applications. Specifically, we test models fine-tuned on OBQA dataset with $4$ domain-specific MMLU subtask ensembles focusing on different professionalities and ARC-E/C datasets to approximate larger and smaller distribution shifts. Metric details are given in Appendix A.6.
5.3 Results
We assess the prediction accuracy and uncertainty calibration of models under both in-distribution and distribution shift scenarios. The in-distribution scenario examines the alignment of the fine-tuned model on the target downstream task, while the distribution shift scenario evaluates the generalizability of the model on novel tasks beyond the fine-tuned domain. These two scenarios combined provides a comprehensive assessment of model robustness in real-world applications where it is essential for the model to excel on its primary task while maintaining the ability to effectively handle unforeseen or out-of-distribution inputs.
Table 2: Performance comparison of different methods fine-tuned on the OBQA dataset with LlaMA2-7B across two smaller distribution shift (DS) tasks and four larger distribution shift tasks. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
ID Smaller DS Larger DS
Metrics Methods OBQA ARC-C ARC-E CS Eng Law Health
ACC $\uparrow$ LoRA $72.1_{0.87}$ $58.6_{1.93}$ $66.5_{3.38}$ $35.5_{2.35}$ $30.8_{1.72}$ $34.9_{1.41}$ $39.1_{1.52}$
MC Drop $74.8_{1.34}$ $58.7_{2.07}$ $66.6_{3.30}$ $36.0_{1.69}$ $30.3_{2.25}$ $35.1_{0.86}$ $39.1_{1.35}$
Ensemble $75.5_{1.4}$ $57.7_{0.78}$ $69.1_{0.48}$ $36.7_{2.18}$ $30.3_{1.13}$ $35.3_{1.02}$ $39.9_{1.89}$
LA $70.8_{1.24}$ $58.7_{0.58}$ $67.9_{0.41}$ $33.7_{1.22}$ $29.6_{1.32}$ $35.4_{0.75}$ $38.5_{1.61}$
MixLoRA $75.5_{2.91}$ $58.5_{1.44}$ $69.2_{1.02}$ $35.2_{2.92}$ $30.3_{0.98}$ $35.9_{0.43}$ $40.6_{1.13}$
UQ4CT $77.3_{1.36}$ $58.8_{1.06}$ $65.8_{1.31}$ $36.2_{1.24}$ $34.1_{2.31}$ $35.8_{1.01}$ $40.0_{1.24}$
ECE $\downarrow$ LoRA $10.2_{1.07}$ $16.7_{2.28}$ $13.3_{2.48}$ $29.7_{2.69}$ $32.3_{1.85}$ $29.2_{3.08}$ $31.0_{2.13}$
MC Drop $10.9_{0.24}$ $16.7_{2.20}$ $13.2_{2.21}$ $23.2_{2.32}$ $31.6_{1.64}$ $28.0_{2.93}$ $25.9_{2.27}$
Ensemble $8.83_{2.35}$ $15.1_{1.09}$ $11.1_{0.99}$ $22.4_{1.32}$ $28.5_{2.13}$ $29.0_{1.37}$ $24.5_{0.39}$
LA $17.2_{1.2}$ $16.2_{0.5}$ $24.4_{0.42}$ $28.6_{2.61}$ $30.5_{1.43}$ $29.5_{1.83}$ $30.7_{1.2}$
MixLoRA $12.9_{1.99}$ $19.0_{1.88}$ $14.5_{2.57}$ $26.4_{3.25}$ $33.7_{1.87}$ $30.3_{2.27}$ $28.3_{1.07}$
UQ4CT $\mathbf{5.0_{1.15}}$ $\mathbf{8.9_{3.46}}$ $\mathbf{6.5_{1.85}}$ $\mathbf{19.6_{2.90}}$ $\mathbf{23.1_{1.17}}$ $\mathbf{25.9_{3.43}}$ $\mathbf{21.9_{3.49}}$
5.3.1 In-distribution Performance
We first evaluate UQ4CT and baseline models fine-tuned on the $4$ common sense reasoning tasks and the climate question answering task under the in-distribution scenario, where models are trained and evaluated on different splits of the same dataset. Note that one of the key advantages of UQ4CT is that uncertainty calibration happens during the fine-tuning stage with little computational overhead, while other UQ methods require repetitive sampling or other post-hoc complexities.
As shown in Table 1, UQ4CT demonstrates notable improvements in uncertainty calibration across a variety of tasks. Across all benchmarks, UQ4CT maintains competitive accuracy (ACC) compared to the baseline methods. For example, on the BoolQ and ClimateQA datasets, UQ4CT achieves accuracy rates of $73.5\%$ and $63.3\%$, respectively. This empirically demonstrates that UQ4CT is capable of maintaining high accuracy with uncertainty calibration, which assures the gain in UQ performance does not compromise accuracy.
The most substantial performance improvement is observed in the reduction of Expected Calibration Error (ECE). UQ4CT consistently outperforms other methods, reducing ECE by more than $25\%$ on average across the evaluated benchmarks. Unlike other methods where the ECE performance is worsened on the more challenging ARC-C benchmark, UQ4CT achieves an ECE score of $6.1$, showcasing the effectiveness of the calibration.
In addition to experiments on LLaMA-2-7B in the main text, we also present additional experiments on fine-tuning Mistral-7B in Appendix A.2 for more comprehensive evaluation of our method. For both LLaMA-2-7B and Mistral-7B models, UQ4CT consistently shows substantial improvements in uncertainty calibration. The improvements are critical in applications where the model’s confidence must align with its predictive accuracy given limited data, particularly in safety-critical and domain-specific tasks.
5.3.2 Performance Under Distribution Shift
Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a). Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios. Similar to the dataset setup in (Yang et al., 2024b), we use ARC-C and ARC-E dataset to simulate smaller distribution shift because the ARC dataset has similar domain focus on general science reasoning, but is generally more challenging and covers a broader range of scientific topics than OBQA. For larger distribution shift, we ensemble the domain-specific MMLU subtasks into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health. These tasks have very broad coverage of the domain task at various knowledge levels ranging from elementary school to professionals. This domain-specificity demonstrates larger distribution shift from OBQA, which is a general common sense reasoning task. Details of the ensemble is provided in Appendix. A.7.
The distribution shift evaluations are provided in Table 2. UQ4CT provides substantial improvements in terms of ECE while maintains similar accuracy for both smaller and larger distribution shift scenarios. For smaller distribution shifts, UQ4CT shows comparable ECE performance as the in-distribution scenario. For the more challenging larger distribution shifts, UQ4CT still achieves the best ECE performance among all baseline models. Note that UQ4CT also achieves competitive prediction accuracy across all domain-specific tasks. This empirically shows that our proposed alignment of the functional epistemic uncertainty with predictive correctness improves generalizability and mitigating the overconfidence problem on the fine-tuned model.
Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintaining similar accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
UQ4CT_Incremental $72.0_{0.19}$ $75.4_{0.81}$ $54.6_{0.95}$ $77.6_{0.43}$ $60.2_{3.17}$
ECE $\downarrow$ UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
UQ4CT_Incremental $4.0_{0.16}$ $9.8_{1.51}$ $13.8_{2.08}$ $10.3_{1.73}$ $12.2_{0.88}$
5.4 Ablation Study
In this section, we conduct ablation studies to investigate the effectiveness of our designed calibration loss, $\mathcal{L}_{cal}$. We first evaluate the incremental weighting performance of the calibration term, which investigates the effectiveness of $\mathcal{L}_{cal}$ at the early stage of fine-tuning. Then, we perform a sensitivity test, where we explore the overall performance impact of $\mathcal{L}_{cal}$. We also conduct an ablation study on the impact of active LoRA experts in Appendix A.4.
5.4.1 Incremental Weighting on Calibration Term
Due to the random initialization of LoRA experts, the predictions during early fine-tuning stage are likely to be incorrect as the model has little knowledge on the functional relationships regarding the data. Thus, it is intuitive to incrementally increase the weight parameter $\beta$ over the calibration term $\mathcal{L}_{\mathrm{cal}}$ in the training loss for the LoRA experts to learn before calibration. We conduct this study by incrementally increase $\beta$ from 00 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
$\beta=\min\left\{1,\frac{\mathrm{current\_grad\_step}}{50}\right\}.$ (15)
We choose $50$ gradient steps from our observation that training loss generally stabilizes after $50$ gradient steps, indicating the LoRA experts have learned some functional relationships from data.
As shown in Table 3, the incremental loss has significantly worse ECE performance across all tasks. This demonstrates the advantage of uncertainty calibration even in the early stage. In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high epistemic uncertainty. Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
Table 4: Performance of UQ4CT with varying $\beta$ values on the OBQA dataset. Prediction accuracy and uncertainty calibration improve with increasing $\beta$, highlighting the effectiveness of $\mathcal{L}_{\mathrm{cal}}$.
$\beta$ ACC $\uparrow$ ECE $\downarrow$
0 $75.5_{2.91}$ $12.9_{1.99}$
0.2 $76.0_{0.6}$ $7.47_{0.78}$
0.5 $75.9_{0.31}$ $7.82_{0.93}$
0.8 $76.6_{0.4}$ $5.94_{1.16}$
1 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
5.4.2 Sensitivity Test on Calibration Term
To further understand the effectiveness of the calibration loss, we perform a sensitivity test of the coefficient $\beta$ in Equation 14. This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities. We evaluate $\beta$ values of $0,0.2,0.5,0.8$ and $1$, where $\beta=0$ resembles the original MixLoRA method.
Results in Table 4 demonstrate the effectiveness of the calibration loss. When $\beta=0$, the model is optimized without calibration on parameter mixtures, resulting in high ECE value. Even with small $\beta=0.2$ or $\beta=0.5$, the ECE scores drastically improved compared to no calibration setting. Finally, when $\beta=1$, the calibration term effectively optimizes the conditional parameter mixtures to generate outputs that fit data distribution well, resulting in lower ECE scores and higher accuracies.
6 Discussion & Conclusion
In this work, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which addresses the overconfidence issues commonly encountered during fine-tuning of large language models. We present a functional perspective on quantifying epistemic uncertainty in LLMs and utilize it for uncertainty-calibrated fine-tuning. By incorporating functional-level epistemic uncertainty quantification with a mixture-of-experts framework, our proposed uncertainty-calibrated training loss effectively addresses the challenge of overconfidence in fine-tuned LLMs by significantly improving uncertainty calibration while maintaining high accuracy. Our evaluations demonstrate that UQ4CT reduces the Expected Calibration Error by more than 25% without compromising accuracy across a variety of downstream tasks, including common-sense and domain-specific reasoning, under in-distribution and out-of-distribution scenarios.
The limitation of UQ4CT lies in its dependency on predictive correctness. For general language modeling tasks such as chat completion, there lacks a clear metric on response correctness. This limits the application of UQ4CT as naively token matching is a poor indicator of semantic correctness due to the ambiguous nature of language. For future work, we are exploring ways to adapt UQ4CT to open-ended problems that lack a definitive optimization objective.
Acknowledgement
This work was supported in part by U. S. Army Research Office under Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, and NSF Grants #2205093, #2146343, #2134274, CDC-RFA-FT-23-0069, DARPA AIE FoundSci and DARPA YFA.
Appendix A Appendix
A.1 Load Balancing Loss
We follow the load balancing loss in [Li et al., 2024]. Given $N$ experts indexed by $i=1$ to $N$ and a batch $B$ with $T$ tokens, the auxiliary loss is computed as:
$\mathcal{L}_{aux}=a\cdot N\cdot\sum^{N}_{i=1}\mathcal{F}_{i}\cdot\mathcal{P}_{i},$ (16)
where
$\mathcal{F}_{i}=\frac{1}{T}\sum_{x\in B}\mathbbm{1}\{argmax_{k}\mathcal{R}(x)_{k}=i\},\mathcal{P}_{i}=\frac{1}{T}\sum_{x\in B}\mathcal{R}(x)_{i}.$ (17)
Here, $\mathcal{R}(\cdot)$ is the top-k router, $\mathcal{F}_{i}$ is the fraction of tokens dispatched to expert $i$ and $\mathcal{P}_{i}$ is the fraction of the router probability allocated for expert $i$. The final loss is multiplied by the expert count $N$ to keep the loss constant as the number of experts varies, and the constant term $a$ is set to $10^{-2}$as a multiplicative coefficient, which is large enough to ensure load balancing while remaining small enough not to overwhelm the primary objective.
A.2 Experimental Results with Mistral-7B
In this section, we present the results using Mistral-7B [Jiang et al., 2023], a different decoder-based LLM backbone. Table 5 shows the results of fine-tuning Mistral-7B on $4$ common-sense reasoning tasks and one domain-specific climate question-answering task.
For each of the tasks, UQ4CT effectively calibrates the parameter mixtures, leading to the best ECE performance in $4$ out of $5$ tasks. This indicates the robustness of UQ4CT across different LLMs.
Table 5: Performance comparison of different methods fine-tuned with Mistral-7B across $4$ common sense reasoning tasks and a domain-specific task. UQ4CT shows significant ECE improvements while maintaining high accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ LoRA $70.3_{0.62}$ $84.8_{0.47}$ $70.2_{0.84}$ $82.8_{0.62}$ $72.5_{1.6}$
MC Drop $69.6_{1.07}$ $84.6_{0.91}$ $69.6_{0.76}$ $82.6_{0.71}$ $72.5_{1.6}$
Ensemble $71.8_{1.29}$ $84.2_{0.66}$ $71.0_{1.41}$ $82.5_{0.6}$ $72.9_{2.88}$
LA $70.7_{1.82}$ $82.4_{2.05}$ $68.5_{3.31}$ $82.5_{0.77}$ $71.6_{1.56}$
MixLoRA $73.1_{0.38}$ $85.5_{1.27}$ $71.2_{1.75}$ $83.3_{1.14}$ $72.0_{1.69}$
UQ4CT $73.6_{0.28}$ $85.9_{0.82}$ $74.4_{0.82}$ $83.7_{1.22}$ $73.2_{1.29}$
ECE $\downarrow$ LoRA $10.17_{0.24}$ $9.46_{1.62}$ $18.42_{1.91}$ $13.3_{0.25}$ $13.72_{2.62}$
MC Drop $10.62_{0.51}$ $8.91_{1.35}$ $18.38_{1.66}$ $13.3_{0.31}$ $13.72_{2.61}$
Ensemble $8.72_{1.13}$ $8.72_{1.49}$ $17.0_{0.97}$ $9.14_{2.82}$ $12.86_{1.78}$
LA $5.33_{2.16}$ $20.3_{5.7}$ $21.27_{4.15}$ $\mathbf{6.41_{3.22}}$ $14.64_{2.21}$
MixLoRA $8.81_{1.03}$ $8.16_{0.99}$ $15.51_{3.86}$ $10.53_{1.73}$ $14.05_{3.09}$
UQ4CT $\mathbf{3.07_{0.83}}$ $\mathbf{5.7_{0.69}}$ $\mathbf{7.04_{0.58}}$ $\underline{7.92_{1.14}}$ $\mathbf{11.4_{1.14}}$
A.3 Incremental Weighting on Calibration Term
Due to the random initialization of LoRA experts, the predictions during early fine-tuning stage are likely to be incorrect as the model has little knowledge on the functional relationships regarding the data. Thus, it is intuitive to incrementally increase the weight parameter $\beta$ over the calibration term $\mathcal{L}_{\mathrm{cal}}$ in the training loss for the LoRA experts to learn before calibration. We conduct this study by incrementally increase $\beta$ from 00 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
$\beta=\min\left\{1,\frac{\mathrm{current\_grad\_step}}{50}\right\}.$ (18)
We choose $50$ gradient steps from our observation that training loss generally stabilizes after $50$ gradient steps, indicating the LoRA experts have learned some functional relationships from data.
As shown in Table 3, the incremental loss has significantly worse ECE performance across all tasks. This demonstrates the advantage of uncertainty calibration even in the early stage. In the beginning, the lack of functional relationships on the training data in LoRA experts lead to high epistemic uncertainty. Thus, UQ4CT encourages exploration over all LoRA experts while UQ4CT_Incremental lacks it due to the small weighting in the beginning.
A.4 Deciding Number of Active Experts
One important aspect of the LoRA MoE architecture is how many experts to activate. Here, we investigate the performance impact of different number of active LoRA experts. We evaluate the model performance with $1$ to $5$ active experts with $8$ in total.
Table 6: Performance comparison of UQ4CT with varying number of experts on OBQA dataset. Top-2 expert selection strategy grants best accuracy and calibration.
Top-K ACC $\uparrow$ ECE $\downarrow$
Top-1 $74.8_{0.62}$ $7.69_{1.96}$
Top-2 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
Top-3 $75.2_{0.8}$ $5.8_{0.81}$
Top-4 $75.8_{0.53}$ $7.67_{0.46}$
Top-5 $75.3_{0.5}$ $6.29_{0.61}$
As shown in Table 6, $2$ active experts give the optimal performance in terms of accuracy and ECE scores. One expert alone cannot capture complicated functional relationships, while more than 2 experts could potentially introduce redundant functional bases to the model, which deviates the output distribution more from data distribution, thus worsening predictive and calibration performance. Additionally, more active experts lead to a more flattened distribution across experts, which hardens the alignment of parameter mixtures during fine-tuning.
A.5 Training Details
We train our model with total of $8$ LoRA experts, and select $2$ experts with the highest probability. For each expert, we use $rank=16$ and $alpha=32$. We use batch size of $16$ to train our model. For climate task, we set the learning rate to $45e-$ 4 and dropout rate to $0.1$ to incorporate the small dataset size. For other tasks, we use $42e-$ 4 as our learning rate with dropout $0.05$. We use AdamW as our optimizer and a cutoff length of $512$ for prompts during training.
The experimental setup for single LoRA based models is similar with LoRA ranks set to $80$ to accommodate the MoE model size. For the ensemble baseline, we use an ensemble size of $8$ with $rank=16$. For Laplace-LoRA, we follow the Laplace hyperparameters in this Github Repository.
A.6 Expected Calibration Error
Expected calibration error (ECE) is a commonly used metric to asses uncertainty quantification performance. ECE measures the alignment between prediction accuracy and model confidence through regrouping the predicted probabilities into $m$ bins. This method then computes the weighted average of the difference between average accuracy and confidence in each bin:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\text{acc}(B_{m})-\text{conf}(B_{m})|,$ (19)
where $|B_{m}|$ is the number of evaluated datapoints in bin $m$, acc and conf is calculated as following:
$\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i}),$ (20)
$\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}P(\hat{y}_{i}).$ (21)
A.7 MMLU Distribution Shift Dataset Composition
  • Computer Science (CS):
    • College Computer Science
    • Computer Security
    • High School Computer Science
    • Machine Learning
  • Engineering (Eng):
    • Electrical Engineering
  • Law:
    • International Law
    • Jurisprudence
    • Professional Law
  • Health:
    • Anatomy
    • Clinical Knowledge
    • College Medicine
    • Human Aging
    • Nutrition
    • Professional Medicine
    • Virology